The decomposition of a generic diatomic element in its standard state is represented by the equation 12X2(g)⟶X(g) Assume that the standard molar Gibbs energy of formation of X(g) is 5.45 kJ·mol−1 at 2000 . K and −59.20 kJ·mol−1 at 3000 . K. Determine the value of the thermodynamic equilibrium constant, ? , at each temperature. At 2000 . K, Δ?f=5.45 kJ·mol−1 . What is ? at that temperature? ? at 2000. K= At 3000 . K, Δ?f=−59.20 kJ·mol−1 . What is K at that temperature? ? at 3000. K=
Answer #1

1) DG0 = -RTlnK

dG0 = 5.45 kj/mol

R = 8.314 j.k-1.mol-1

T = 2000 k

K = equilibrium constant = ?

5.45*10^3 = -8.314*2000lnK

K = 0.720

2) DG0 = -RTlnK

dG0 = -59.20 kj/mol

R = 8.314 j.k-1.mol-1

T = 3000 k

K = equilibrium constant = ?

-59.20*10^3 = -8.314*3000lnK

K = 10.735